What is the domain of f(x) = 10/x^2-2x ?
Answer by stanbon(75887)

What is the domain of f(x) = 10/x^2-2x
---
The denominator cannot be zero.
Solve x^2-2x = 0
x(x-2) = 0
x = 0 or x = 2
-----
Domain is All Real Numbers except x = 0 and x = 2
=======================================================
Cheers,
Stan H.